# MetaBox-v2: A Unified Benchmark Platform for Meta-Black-Box Optimization

### v1

@@ Abstract
Meta-Black-Box Optimization (MetaBBO) streamlines the automation of optimization algorithm design through meta-learning. It typically employs a bi-level structure: the meta-level policy undergoes meta-training to reduce the manual effort required in developing algorithms for low-level optimization tasks. The original MetaBox (2023) provided the first open-source framework for reinforcement learning-based single-objective MetaBBO. However, its relatively narrow scope no longer keep pace with the swift advancement in this field. In this paper, we introduce MetaBox-v2 (https://github.com/MetaEvo/MetaBox) as a milestone upgrade with four novel features: 1) a unified architecture supporting RL, evolutionary, and gradient-based approaches, by which we reproduce $23$ up-to-date baselines; 2) efficient parallelization schemes, which reduce the training/testing time by $10-40$x; 3) a comprehensive benchmark suite of $18$ synthetic/realistic tasks ($1900$+ instances) spanning single-objective, multi-objective, multi-model, and multi-task optimization scenarios; 4) plentiful and extensible interfaces for custom analysis/visualization and integrating to external optimization tools/benchmarks. To show the utility of MetaBox-v2, we carry out a systematic case study that evaluates the built-in baselines in terms of the optimization performance, generalization ability and learning efficiency. Valuable insights are concluded from thorough and detailed analysis for practitioners and those new to the field.
@@ 1 Introduction
2) Efficient Parallelization: MetaBBO approaches are typically time-consuming due to the nested bi-level structure. In MetaBox-v2, we introduce two parallel schemes: vectorized optimization environment and instance-level distributed evaluation, to accelerate MetaBBO by $10-40$x.
@@ 2 Related Works
MetaBBO. We first illustrate the bi-level paradigm of Meta-Black-Box-Optimization (MetaBBO) [6] in Figure 2. In low-level optimization environment, a BBO optimizer $\mathcal{A}$ is maintained to optimize a problem $p$ sampled from distribution $\mathcal{P}$. At each optimization step $t$, optimization status features are extracted from the current optimization process (such as population and objective values information). Then in meta level, an algorithm design policy $\pi_{\theta}$ (with learnable parameters $\theta$) outputs a desired design $\omega_{i}^{t}$ by $\omega_{i}^{t}=\pi_{\theta}(s_{i}^{t})$. $\mathcal{A}$ optimizes $p$ by $\omega_{i}^{t}$ for one step. A performance measurement function $r_{t}$ is used to evaluate the performance gain obtained by this algorithm design decision. Suppose $T$ optimization steps are allowed for the low-level optimization process, then $\pi_{\theta}$ is meta-trained to maximize a meta-objective formulated as: $J(\theta)=\mathbb{E}_{p\in\mathcal{P}}[\sum_{t=1}^{T}r_{t}]$, which is expectation of accumulated single step performance gain over all problem instances in $\mathcal{P}$. In practice, a training problem set serves as the distribution $\mathcal{P}$.
@@
Following this paradigm, a wide array of MetaBBO approaches have been proposed, which further diverge into four same-end branches according to the learning techniques they adopt [6, 7]. The four branches are: 1) MetaBBO-RL: those first model the algorithm design process as a Markov Decision Process (MDP), then employ effective RL techniques to learn well-performing policies. Initial works such as DEDDQN [27], LDE [28], DEDQN [29] and RLEPSO [30] focus on designing dynamic configuration strategy for BBO optimizer. Following them, in-depth exploration include high-capacity neural policy [31, 32], complete optimizer generation [33, 34], optimization feature learning [35, 36] and efficient offline learning [37, 38]; 2) MetaBBO-SL: this branch originates from RNN-opt [11], where given a solution as input, a RNN is used to auto-regressively iterate it for better solution. The RNN is trained by minimizing the differentiable objective function. Although this paradigm requires white-box (differentiable) problems for training, recent works such as GLHF [13] and B2Opt [12] demonstrate that only training on synthetic problems is sufficient for generalization towards unseen problems; 3) MetaBBO-NE: where a the meta-level policy is learned by evolutionary optimization on its net parameters [15, 14, 39]; 4) MetaBBO-ICL: where a general LLM serves as either the low-level optimizer [17, 18, 19], or a meta-level configuration policy [40, 41]. The textual optimization process information is regarded as the context and learned by the LLM to propose algorithm designs.
@@
MetaBox-v2 5 RL,SL, NE,ICL 541+1393 23 rich
Related Benchmarks. A comparison of MetaBox-v2 to representative and up-to-date BBO benchmarks is presented in Table 1 to show the novelty of our work. Apart from those have been reviewed and compared in MetaBox [8], latest efforts on developing BBO benchmark include: 1) Engineering [47]: a collection of real world engineering optimization problems such as heat exchanger network design, industrial chemical process optimization, etc; 2) MA-BBOB [48]: a many-affine problem sets constructed from COCO [42] by interpolation operations on COCO’s problem instances, resulting in diversified synthetic instances; 3) BBOPlace [49]: chip placement tasks which represents complex and challenging optimization scenarios; 4) PyPop7 [50]: a comprehensive benchmark platform featured by massive number of BBO optimizers, which include decades of advanced optimizers with different types and specialized scenarios. 5) EvoX [51]: a high-efficiency benchmark platform featured by its distributed GPU-accelerated evaluation framework, with over $100$x speedups than traditional BBO benchmarks such as CoCo and PyPop7. Notably, MetaBox, across its two versions, remains the sole platform that supports MetaBBO’s bi-level framework to streamline the development and benchmarking processes in this research domain.
@@ 3.1 Unified MetaBBO Interface
bbob-10D[42] SOO 10D 2E4 24 synthetic Single-objective instances in CoCo BSD-3-Clause
bbob-30D[42] SOO 30D 5E4 24 synthetic Single-objective instances in CoCo BSD-3-Clause
bbob-noisy-10D[42] SOO 10D 2E4 24 synthetic bbob-10D with gaussian noise BSD-3-Clause
bbob-noisy-30D[42] SOO 30D 5E4 24 synthetic bbob-30D with gaussian noise BSD-3-Clause
hpo-b[63] SOO 2-16D 2E3 935 realistic Hyper-parameter optimization MIT License
@@
protein[64] SOO 12D 2E3 280 realistic Simplified protein-docking instances Attribution 4.0
@@
ne[51] LSGO $\geq$1000D 2.5E3 66 realistic Neuroevolution for control tasks GPL-3.0
@@
uf[58] MOO 30D 5E3 10 synthetic Multi-objective problem instances Apache-2.0
@@
wfg[59] MOO 12-38D 5E3 117 synthetic Complex multi-objective problems Apache-2.0
moo-uav[56] MOO 30D 2.5E3 56 realistic Multi-objective form of uav Apache-2.0
mmo[61] MMO 1-20D 5E4-4E5 20 synthetic Standard multi-modal problems Simplified BSD
@@ 3.2 Efficiency Optimization
Parallel Testing. MetaBox-v2 provides Ray-based parallel scheme [69] for distributed testing of MetaBBO/BBO baselines. As illustrated in the right of Figure 3, given a MetaBBO’s meta-level agent and a testsuite, we first copy the agent for each testing run and use Ray to construct the corresponding sub-tasks. Then all sub-tasks are distributed into independent CPU/GPU cores for parallel testing. The testing results in these sub-tasks are aggregated automatically by Ray’s handler. By decomposing parallelism into two orthogonal dimensions: (a) distributed solving across problem instances, and (b) parallel execution of independent test runs, we introduce four parallel modes: Mode-1: partially enable (a) by distributing batched problem instances; Mode-2: disable (a) and enable (b) for test run distribution; Mode-3: enable (a) and disable (b) for instance-wise solving; and Mode-4: enable (a) and (b) for full parallelization. This design enables flexible hardware scaling.
@@ 3.3 Novel Evaluation Metrics
[…]
@@
Anti-NFL Indicator. Recall that the core motivation of MetaBBO is to learn generalizable policy towards unseen problems. This is somewhat against the well-known no-free-lunch theorem [5]. We hence propose a novel indicator named as $\operatorname{Anti-NFL}$, which could reflect the performance variance of a given algorithm $\mathcal{A}$ on unseen testsuites apart from the one it was trained. Suppose $\mathcal{A}$ is trained on $\mathbb{D}_{\text{train}}$, and tested on other $B$ testsuites: $\{\mathbb{D}_{\text{test}}^{(b)}\}_{b=1}^{B}$. After obtaining all of the metadata by testing the final model $\mathcal{A}^{(G)}$ on these testsuites, the $\operatorname{Anti-NFL}$ is computed as:
@@ 4.2 Platform’s Acceleration Performance (RQ1)
We use vectorized environment with batch_size as $16$ to accelerate the training of all involved MetaBBO baselines. Due to the space limitation, we selectively illustrate in Figure 4 the performance improvement curves of $8$ baselines, where y-axis denotes normalized performance on testing set of bbob-10D. Compared to original MetaBox, MetaBox-v2 consistently accelerates MetaBBO baselines by at most $10$x. We can observe that the concrete acceleration may varies on different baselines, this is because the differences of the internal logic and communication cost among the baselines.
We also illustrate the acceleration performance of MetaBox-v2 compared to original MetaBox in Figure 5 in terms of testing efficiency, where y-axis denotes the throughput of evaluation process measured by the number of instance test runs per second. We compare the throughput of the $4$ Ray modes in MetaBox-v2 to original MetaBox, and the results show that even the simplest distribution Mode-1 could significantly accelerate the testing workflow. If users have advanced hardware, the distribution Mode-4 could introduce no less than $40$x acceleration.
@@ 4.3 Generalization Performance Comparisons among Baselines (RQ2)
Table 3: In-distribution optimization performances of baselines over bbob-10D, with gray box labeling the best. Due to the space limitation, results for $8$ of $16$ problem instances in bbob-10D’s testing set are presented here while the complete results can be accessed at this online page.
@@ 5 Discussion
Takeaways. This work proposes MetaBox-v2 as a milestone upgrade for its predecessor, introducing several key architecture adjustments that establish a comprehensive benchmark platform for the MetaBBO research. The fundemental advancements not only enable unified development and evaluation for various MetaBBO paradigms and diverse optimization problem types; but also support streamlined parallel acceleration for training/evaluation by $10$x-$40$x. With the expanded baseline library ($8\rightarrow 23$) and testsuite library ($3\rightarrow 18$, $1900+$ problem instances), we conduct a rigorous case study, which discloses insights including but not limited to: 1) Current literature overestimates MetaBBO capabilities through narrow evaluation practices, with MetaBox-v2 exposing substantial performance gaps in cross-domain settings. 2) Out-of-distribution generalization demands special attention, since our analysis reveals that overfitting persists across baseline algorithms. 3) Effective MetaBBO assessment requires multidimensional analysis (optimization efficacy, learning efficiency, generalization robustness) beyond traditional convergence metrics.

### v2

@@ Abstract
Meta-Black-Box Optimization (MetaBBO) streamlines the automation of optimization algorithm design through meta-learning. It typically employs a bi-level structure: the meta-level policy undergoes meta-training to reduce the manual effort required in developing algorithms for low-level optimization tasks. The original MetaBox (2023) provided the first open-source framework for reinforcement learning-based single-objective MetaBBO. However, its relatively narrow scope no longer keep pace with the swift advancement in this field. In this paper, we introduce MetaBox-v2 (https://github.com/MetaEvo/MetaBox) as a milestone upgrade with four novel features: 1) a unified architecture supporting RL, evolutionary, and gradient-based approaches, by which we reproduce $23$ up-to-date baselines; 2) efficient parallelization schemes, which reduce the training/testing time by 104010-40x; 3) a comprehensive benchmark suite of $18$ synthetic/realistic tasks (19001900+ instances) spanning single-objective, multi-objective, multi-model, and multi-task optimization scenarios; 4) plentiful and extensible interfaces for custom analysis/visualization and integrating to external optimization tools/benchmarks. To show the utility of MetaBox-v2, we carry out a systematic case study that evaluates the built-in baselines in terms of the optimization performance, generalization ability and learning efficiency. Valuable insights are concluded from thorough and detailed analysis for practitioners and those new to the field.
@@ 1 Introduction
2) Efficient Parallelization: MetaBBO approaches are typically time-consuming due to the nested bi-level structure. In MetaBox-v2, we introduce two parallel schemes: vectorized optimization environment and instance-level distributed evaluation, to accelerate MetaBBO by 104010-40x.
@@ 2 Related Works
MetaBBO. We first illustrate the bi-level paradigm of Meta-Black-Box-Optimization (MetaBBO) [6] in Figure 2. In low-level optimization environment, a BBO optimizer $\mathcal{A}$ is maintained to optimize a problem $p$ sampled from distribution $\mathcal{P}$. At each optimization step $t$, optimization status features are extracted from the current optimization process (such as population and objective values information). Then in meta level, an algorithm design policy $\pi_{\theta}$ (with learnable parameters $\theta$) outputs a desired design $\omega_{i}^{t}$ by $\omega_{i}^{t}=\pi_{\theta}(s_{i}^{t})$. $\mathcal{A}$ optimizes $p$ by $\omega_{i}^{t}$ for one step. A performance measurement function $r_{t}$ is used to evaluate the performance gain obtained by this algorithm design decision. Suppose $T$ optimization steps are allowed for the low-level optimization process, then $\pi_{\theta}$ is meta-trained to maximize a meta-objective formulated as: $J(\theta)=\mathbb{E}_{p\in\mathcal{P}}[\sum_{t=1}^{T}r_{t}]$, which is expectation of accumulated single step performance gain over all problem instances in $\mathcal{P}$. In practice, a training problem set serves as the distribution $\mathcal{P}$. It is important to note that in this brief introduction we use $\omega$ as an abstract notation for algorithm design, considering that MetaBBO provides a universal concept framework for diverse algorithm design tasks such as algorithm selection, algorithm configuration etc. $\omega$ is hence a flexible and abstract concept to represent specific design choice under the given context. For a comprehensive summary of different algorithm design categories, we refer to MetaBBO survey [6] for further reading.
@@
Following this paradigm, a wide array of MetaBBO approaches have been proposed, which further diverge into four same-end branches according to the learning techniques they adopt [6, 7]. The four branches are: 1) MetaBBO-RL: those first model the algorithm design process as a Markov Decision Process (MDP), then employ effective RL techniques to learn well-performing policies. Initial works such as DEDDQN [27], LDE [28], DEDQN [29] and RLEPSO [30] focus on designing dynamic configuration strategy for BBO optimizer. Following them, in-depth exploration include high-capacity neural policy [31, 32], complete optimizer generation [33, 34], optimization feature learning [35, 36] and efficient offline learning [37, 38]; 2) MetaBBO-SL: this branch originates from RNN-opt [11], where given a solution as input, a RNN is used to auto-regressively iterate it for better solution. The RNN is trained by minimizing the differentiable objective function. Although this paradigm requires white-box (differentiable) problems for training, recent works such as GLHF [13] and B2Opt [12] demonstrate that only training on synthetic problems is sufficient for generalization towards unseen problems; 3) MetaBBO-NE: where a the meta-level policy is learned by evolutionary optimization on its net parameters [15, 14, 39]; 4) MetaBBO-ICL: where a general LLM serves as either the low-level optimizer [17, 18, 19], or a meta-level configuration policy [40, 41]. The textual optimization process information is regarded as the context and learned by the LLM to propose algorithm designs. The fast development in MetaBBO field anticipates corresponding benchmarks.
@@
MetaBox-v2 5
RL,SL,
NE,ICL
541+1393 23 rich
Related Benchmarks. A comparison of MetaBox-v2 to representative and up-to-date BBO benchmarks is presented in Table 1 to show the novelty of our work. Apart from those have been reviewed and compared in MetaBox [8], latest efforts on developing BBO benchmark include: 1) Engineering [47]: a collection of real world engineering optimization problems such as heat exchanger network design, industrial chemical process optimization, etc; 2) MA-BBOB [48]: a many-affine problem sets constructed from COCO [42] by interpolation operations on COCO’s problem instances, resulting in diversified synthetic instances; 3) BBOPlace [49]: chip placement tasks which represents complex and challenging optimization scenarios; 4) PyPop7 [50]: a comprehensive benchmark platform featured by massive number of BBO optimizers, which include decades of advanced optimizers with different types and specialized scenarios. 5) EvoX [51]: a high-efficiency benchmark platform featured by its distributed GPU-accelerated evaluation framework, with over 100100x speedups than traditional BBO benchmarks such as CoCo and PyPop7. Notably, MetaBox, across its two versions, remains the sole platform that supports MetaBBO’s bi-level framework to streamline the development and benchmarking processes in this research domain.
@@ 3.1 Unified MetaBBO Interface
bbob-10D[42] SOO 10D 2E4 24 synthetic
Single-objective instances in CoCo
BSD-3-Clause
bbob-30D[42] SOO 30D 5E4 24 synthetic
Single-objective instances in CoCo
BSD-3-Clause
bbob-noisy-10D[42] SOO 10D 2E4 24 synthetic
bbob-10D with gaussian noise
BSD-3-Clause
bbob-noisy-30D[42] SOO 30D 5E4 24 synthetic
bbob-30D with gaussian noise
BSD-3-Clause
hpo-b[63] SOO 2-16D 2E3 935 realistic
Hyper-parameter optimization
MIT License
@@
protein[64] SOO 12D 2E3 280 realistic
Simplified protein-docking instances
Attribution 4.0
@@
ne[51] LSGO $\geq$1000D 2.5E3 66 realistic
Neuroevolution for control tasks
GPL-3.0
@@
uf[58] MOO 30D 5E3 10 synthetic
Multi-objective problem instances
Apache-2.0
@@
wfg[59] MOO 12-38D 5E3 117 synthetic
Complex multi-objective problems
Apache-2.0
moo-uav[56] MOO 30D 2.5E3 56 realistic
Multi-objective form of uav
Apache-2.0
mmo[61] MMO 1-20D 5E4-4E5 20 synthetic
Standard multi-modal problems
Simplified BSD
@@ 3.2 Efficiency Optimization
Parallel Testing. MetaBox-v2 provides Ray-based parallel scheme [69] for distributed testing of MetaBBO/BBO baselines. As illustrated in the right of Figure 3, given a MetaBBO’s meta-level agent and a testsuite, we first copy the agent for each testing run and use Ray to construct the corresponding sub-tasks. Then all sub-tasks are distributed into independent CPU/GPU cores for parallel testing. The testing results in these sub-tasks are aggregated automatically by Ray’s handler. To show the detail, we first assume a testing scenario where $N$ problem instances compose the testing set, $B$ baselines to be tested, and $R$ independent runs to ensure the statistical robustness. Naive testing procedure requires nested 3-layer loop to complete all these $N\times B\times R$ testing runs. In contrast, in MetaBox-v2, we provide four parallel testing modes to achieve fine-grained testing efficiency optimization: 1) mode-1: N cpu cores are used to distribute the N testing problem instances, on each core, line2-line3 are executed as two-layer loop; 2) mode-2: R cpu cores are used to distribute the R independent runs, while on each core, line1-line2 are executed as two-layer loop; 3) mode-3: NxB cpu cores are used to distribute the N problem instances and B baselines, while on each core, line3 are executed as one-layer loop; 4) mode-4: NxBxR cpu cores are used to distribute all evaluation tasks, hence there is no loop anymore. The hardware requirement from mode-1 to mode-4 is incremental. By decomposing parallelism into two orthogonal dimensions as above: (a) distributed solving across problem instances, and (b) parallel execution of independent test runs, we provide sufficient flexibility for users to accelerate their programs based on their specific hardware conditions.
@@ 3.3 Novel Evaluation Metrics
where "X" is a list of each generation’s solution population, "Y" is a list of the objective values, "T" is the wall-clock time consumed for optimizing $p_{i}$. Then the overall metadata $md(\mathcal{A},\mathbb{D})$ is aggregated from each $p_{i}$: $\{\texttt{"problem\_type":SOO},\ \ \texttt{"all\_data":}[md_{1},...,md_{N}]\}$, where "problem_type" is the optimization types of $\mathbb{D}$. Notably, $md(\mathcal{A},\mathbb{D})$ provide significant convenience for computing various metrics. For example, a normalized performance metric widely used in existing MetaBBO approaches [15, 31, 70, 13] can be easily computed as $\operatorname{Perf}(\mathcal{A},\mathbb{D})=\frac{1}{N\times K}[\sum_{i=1}^{N}\sum_{j=1}^{K}\frac{Y_{i,j}^{*}-p_{i}^{*}}{Y_{i,j}^{0}-p_{i}^{*}}]$, where $Y_{i,j}^{0}$ and $Y_{i,j}^{*}$ are the initial and final objective values, $p_{i}^{*}$ is the optimal. We next showcase two customized metrics based on the metadata.
@@
Anti-NFL Indicator. Recall that the core motivation of MetaBBO is to learn generalizable policy towards unseen problems. This is somewhat against the well-known no-free-lunch (NFL) theorem [5]. We hence propose a novel indicator named as $\operatorname{Anti-NFL}$, which could reflect the performance variance of a given algorithm $\mathcal{A}$ on unseen testsuites apart from the one it was trained. Suppose $\mathcal{A}$ is trained on $\mathbb{D}_{\text{train}}$, and tested on other $B$ testsuites: $\{\mathbb{D}_{\text{test}}^{(b)}\}_{b=1}^{B}$. After obtaining all of the metadata by testing the final model $\mathcal{A}^{(G)}$ on these testsuites, the $\operatorname{Anti-NFL}$ is computed as:
@@ 4.2 Platform’s Acceleration Performance (RQ1)
We use vectorized environment with batch_size as $16$ to accelerate the training of all involved MetaBBO baselines. Due to the space limitation, we selectively illustrate in Figure 4 the performance improvement curves of $8$ baselines, where y-axis denotes normalized performance on testing set of bbob-10D. Compared to original MetaBox, MetaBox-v2 consistently accelerates MetaBBO baselines by at most 1010x. We can observe that the concrete acceleration may varies on different baselines, this is because the differences of the internal logic and communication cost among the baselines. An important note is that the irregular record points in Figure 4 is due to the unstable multi-processing. For parallel cases, we draw 5 points (every 20 training epochs), which might present irregular patterns since the x-axis denotes the training time.
We also illustrate the acceleration performance of MetaBox-v2 compared to original MetaBox in Figure 5 in terms of testing efficiency, where y-axis denotes the throughput of evaluation process measured by the number of instance test runs per second. We compare the throughput of the $4$ Ray modes in MetaBox-v2 to original MetaBox, and the results show that even the simplest distribution Mode-1 could significantly accelerate the testing workflow. If users have advanced hardware, the distribution Mode-4 could introduce no less than 4040x acceleration.
@@ 4.3 Generalization Performance Comparisons among Baselines (RQ2)
Table 3: In-distribution optimization performances of baselines over bbob-10D, with gray box labeling the best. Due to the space limitation, results for 88 of $16$ problem instances in bbob-10D’s testing set are presented here while the complete results can be accessed at this online page.
@@ 5 Discussion
Takeaways. This work proposes MetaBox-v2 as a milestone upgrade for its predecessor, introducing several key architecture adjustments that establish a comprehensive benchmark platform for the MetaBBO research. The fundemental advancements not only enable unified development and evaluation for various MetaBBO paradigms and diverse optimization problem types; but also support streamlined parallel acceleration for training/evaluation by 1010x-4040x. With the expanded baseline library ($8\rightarrow 23$) and testsuite library ($3\rightarrow 18$, $1900+$ problem instances), we conduct a rigorous case study, which discloses insights including but not limited to: 1) Current literature overestimates MetaBBO capabilities through narrow evaluation practices, with MetaBox-v2 exposing substantial performance gaps in cross-domain settings. 2) Out-of-distribution generalization demands special attention, since our analysis reveals that overfitting persists across baseline algorithms. 3) Effective MetaBBO assessment requires multidimensional analysis (optimization efficacy, learning efficiency, generalization robustness) beyond traditional convergence metrics.
@@
Acknowledgments and Disclosure of Funding
This work was supported in part by National Natural Science Foundation of China (Grant No. 62276100), in part by Guangzhou Science and Technology Elite Talent Leading Program for Basic and Applied Basic Research (Grant No. SL2024A04J01361), in part by the Guangdong Provincial Natural Science Foundation for Outstanding Youth Team Project (Grant No. 2024B1515040010), and in part by the Fundamental Research Funds for the Central Universities (Grant No. 2025ZYGXZR027).